the function f : x ı→ a + b cos x, is defined for 0 ≤ x ≤ 2π. Given that f(0) = 10 and that f(2π/3) = 1 , find (i) the values of a and b, (ii) the range of f, (iii) the exact value of f(5π/6)
1
Expert's answer
2020-06-15T19:00:08-0400

I. "f(x)=a+b cos x, 0\le x \le 2\pi."

"f(0)=10," hence "a+b cos 0=10, a+b\cdot 1=10, a+b=10."

"f(\frac{2\pi}{3})=1," hence "a+b cos \frac{2\pi}{3}=1, a+b\cdot (-\frac{1}{2})=1," when multiply this equation by 2, we have "2a-b=2."

We have the system of linear equations "\begin{cases} a+b=10, \\ 2a-b=2. \end{cases}"

If we add the equations, we have "3a=12," hence "a=4."

Then "4+b=10, b=6."

Answer. "a=4, b=6."

II. "f(x)=4+6 cos x."

The magnitude of the function is 6.

The lower bound of the range for the function is found when "cos x=-1," and it's equal to "y=4+6\cdot (-1)=-2."

The upper bound of the range for the function is found when "cosx=1", and it's equal to "y=4+6\cdot 1=10."

Hence the range of the f is "-2 \le y \le 10," interval notation "[-2, 10]."

Answer. "[-2, 10]."

III. "f(\frac{5\pi}{6})=4+6cos\frac{5\pi}{6}=4+6\cdot (-\frac{\sqrt{3}}{2})=4-3\sqrt{3}."

Answer. "4-3\sqrt{3}."